# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v3

@@ III-D Local Keypoint Feature Aggregator
Next, for each keypoint, we construct a local neighborhood in both coordinate and feature spaces using the $K$-nearest neighbor ($K\!N\!N$) algorithm. This yields coordinate neighborhoods $\mathcal{P}_{K}\in\mathbb{R}^{N_{kpt}\times K\times 3}$ and feature neighborhoods $\mathcal{F}_{K}\in\mathbb{R}^{N_{kpt}\times K\times d}$, where $K$ is the neighborhood size. We then concatenate the spatial and feature descriptors and apply standard normalization to obtain fused local group features $\mathcal{F}_{K}^{\prime}\in\mathbb{R}^{N_{kpt}\times K\times(d+3)}$, which jointly encode geometric structure and semantic information:
@@
To further enhance the representational capacity of each keypoint, we concatenate the normalized neighbor features $\mathcal{F}^{\prime}_{K}$ with the corresponding central keypoint feature $\mathcal{F}_{kpt}$, resulting in expanded representations $\mathcal{F}^{\prime\prime}_{K}\in\mathbb{R}^{N_{kpt}\times K\times(2\times d+3)}$. This operation allows each keypoint to incorporate both its own feature and contextual information from its spatial neighborhood. We then apply an affine transformation to $\mathcal{F}^{\prime\prime}_{K}$, parameterized by learnable weights $\alpha$ and $\beta$, to capture potential rigid transformations in local geometry and adaptively refine the feature distribution. The transformed feature set ^$\hat{\mathcal{F}}_{K}\in\mathbb{R}^{N_{kpt}\times K\times(2\times d+3)}$ is computed as follows:
@@ III-E Global Keypoint Feature Aggregator
As illustrated in Fig. 3(b), GKFA employs a bidirectional Mamba-based architecture [26] to efficiently model long-range dependencies and capture object-level structural information. Compared to traditional Transformers, which suffer from quadratic complexity with respect to sequence length, Mamba operates in linear time, making it a more scalable and efficient choice for processing keypoint sequences in point clouds.
To fully exploit Mamba’s bidirectional modeling capabilities, it is necessary to construct both forward and backward feature sequences. A common practice in prior works is to reverse the input sequence order, a technique we refer to as Point Sequence Flipping (PSF). However, point cloud data lack an inherent spatial order, and reversing the index sequence may place geometrically unrelated keypoints next to each other, thereby disrupting the continuity of feature learning. To address this issue, we propose a simple yet effective strategy termed Feature Sequence Flipping (FSF). Rather than reordering keypoints, FSF reverses the feature channels within each keypoint vector while preserving the original spatial arrangement of the point cloud. Formally, given a keypoint feature vector $\mathcal{F}_{local}(i)\in\mathbb{R}^{d}$ expressed as $[f_{1}(i),f_{2}(i),\ldots,f_{d}(i)]$, we define FSF as:
$\text{FSF}(\mathcal{F}_{local}(i))=[f_{d}(i),f_{d-1}(i),\ldots,f_{1}(i)].$ (9)
@@
$\mathcal{F}_{global}=\text{SSM}(\mathcal{F}_{local},\text{FSF}(\mathcal{F}_{local}))+\mathcal{F}_{local}.$ (10)
@@
To ensure that the aggregated keypoints retain meaningful and expressive geometric structure, we utilize the reconstruction-based supervision mechanism [20]. Specifically, we employ an MLP layer to regress dense positional offsets from the global keypoint features. Given $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$, we repeat each feature vector $m=\lfloor N_{rec}/N_{kpt}\rfloor$ times to construct an expanded representation $\mathcal{F}_{global}^{\prime}\in\mathbb{R}^{N_{rec}\times d}$, where $N_{rec}$ is the predefined number of reconstruction points (set to 960 in our experiments). The MLP predicts a set of 3D offsets $\mathcal{O}_{\mathcal{P}}\in\mathbb{R}^{N_{rec}\times 3}$:
@@ III-F Pose and Size Estimator
After obtaining semantically consistent and geometrically informative keypoints and their associated features, we adopt a standard Pose and Size Estimator, as in previous work [14, 20], to infer the object’s 6D pose and size. Specifically, we first transform the predicted keypoint coordinates into the NOCS space using the ground-truth rotation $\textit{R}_{gt}$, translation $\textit{t}_{gt}$, and scale $\textit{s}_{gt}$:
@@
We then predict the NOCS keypoints from the extracted features via a self-attention mechanism followed by an MLP layer:
@@
[…]
@@ IV-A1 Datasets
Following previous works [20, 21], we evaluate the effectiveness of the proposed INKL-Pose on three widely used benchmarks: CAMERA25 [10], REAL275 [10], and the more challenging HouseCat6D [31] dataset. CAMERA25 is a large-scale synthetic dataset covering six object categories, containing 300k synthetic RGB-D images with objects rendered on virtual backgrounds, of which 25k images are used for evaluation. REAL275 serves as a real-world counterpart to CAMERA25, covering the same six categories. The training set comprises 4.3k images captured from 7 scenes, while the test set includes 2.75k images from 6 different scenes, with three distinct object instances per category. HouseCat6D is a more complex, real-world multimodal dataset featuring 10 object categories. The training set contains 20k RGB-D images collected across 34 indoor scenes, and the test set includes 3k images from 5 additional scenes. Notably, this dataset poses greater challenges due to its marker-free capture process, dense clutter, and severe occlusion.
@@ IV-A2 Evaluation Metrics
Following previous works [16, 20, 21], we evaluate the performance of INKL-Pose using two widely adopted metrics.
@@
  • $n$°$m$cm. This metric evaluates the 6D pose accuracy by jointly considering rotation and translation errors. A prediction is regarded as correct if the rotation error is less than $n$° and the translation error is below $m$ cm. Following standard protocol, we report results under four commonly used thresholds: 5°2cm, 5°5cm, 10°2cm, and 10°5cm across all three datasets.
@@ IV-A3 Implementation Details
We utilize the off-the-shelf Mask R-CNN [32] to generate instance masks and resize all input images to 224$\times$224 pixels. For each corresponding object instance, we uniformly sample 1,024 points from the segmented point cloud. To enhance robustness, we apply data augmentation strategies consistent with prior works [20, 14], including random rotations (within 0–20 degrees), translations (in the range of –0.02 to 0.02 meters), and scalings (from 0.8 to 1.2). For network configuration, the feature dimensions are set as follows: $d_{1}$ = 128, $d_{2}$ = 128 , $d_{3}$ = 64 , and the fused feature dimension is $d=256$. In the local feature aggregation stage, the neighborhood size $K$ is set to 4. Additionally, the entire feature aggregation module consists of $S$ = 12 stacked stages, each comprising both local and global aggregation components to progressively refine the keypoint features. In the keypoint generation stage, the network predicts 96 keypoints per object instance. For keypoint supervision, we sample $N_{fps}$ = 120 reference points to ensure sufficient surface coverage and spatial alignment with the predicted keypoints. In the separation loss, the number of nearest neighbors is set to $M$ = 2, which provides a lightweight yet effective constraint that mitigates keypoint clustering while preserving flexibility in spatial distribution. In the dense point cloud reconstruction stage, each keypoint is repeated 10 times, yielding $N_{rec}$ = 960 reconstructed points. This setting closely matches the resolution of the input point cloud (1,024 points), providing a stable foundation for reconstruction supervision without incurring excessive computational cost.
Regarding optimization, the weights of the loss terms are set as {$w_{sep},w_{surf},w_{sim},w_{map},w_{pose}\}=\{10.0,10.0,15.0,2.0,0.3\}$. The model is trained using the Adam optimizer [33] with a cyclical learning rate schedule [34], where the learning rate varies between $2\times 10^{-5}$ and $5\times 10^{-4}$. All experiments are conducted on a single NVIDIA RTX 4090D GPU with a batch size of 64. For CAMERA25 and REAL275, we train the model using both synthetic and real-world data with a synthetic-to-real ratio of 3:1. For HouseCat6D, the model is trained exclusively on its corresponding dataset.
@@
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$ (M) $\downarrow$
SPD [ECCV’20][11] 77.3 53.2 19.3 21.4 43.2 54.1 18.3
SGPA [ICCV’21][13] 80.1 61.9 35.9 39.6 61.3 70.7 23.3
GPV-Pose [CVPR’22][35] 83.0 64.4 32.0 42.9 55.0 73.3 8.6
DPDN [ECCV’22][14] 83.4 76.0 46.0 50.7 70.4 78.4 24.6
IST-Net [ICCV’23][17] 82.5 76.6 47.5 53.4 72.1 80.5 26.8
Query6DoF [ICCV’23][15] 82.5 76.1 49.0 58.9 68.7 83.0 19.7
VI-Net [ICCV’23][18] - - 50.0 57.6 70.8 82.1 28.9
CatFormer [AAAI’24][12] 83.1 73.8 47.7 53.7 69.0 79.5 -
CLIPose [TCSVT’24][16] - - 48.7 58.3 70.4 85.2 -
SecondPose [CVPR’24][19] - - 56.2 63.6 74.7 86.0 38.1
AG-Pose [CVPR’24][20] 83.8 77.6 56.2 62.3 73.4 81.2 28.9
AG-Pose (DINOv2) [CVPR’24][20] 84.1 80.1 57.0 64.6 75.1 84.7 11.0
SpherePose [ICLR’25][21] 84.0 79.0 58.2 67.4 76.2 88.2 -
KeyPose [AAAI’25][36] 84.2 80.0 57.7 66.0 78.8 88.0 -
GCE-Pose [CVPR’25][37] 84.1 79.8 57.0 65.1 75.6 86.3 -
Ours 84.0 80.6 63.2 68.3 80.9 88.0 16.7
TABLE I: Quantitative comparisons with state-of-the-art methods on the REAL275 dataset[10]. The best results are in bold. The symbol indicates that only the trainable parameter count is reported. The symbol -- indicates that the data is not available or not reported.
@@
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$ (M) $\downarrow$
SPD [ECCV’20] [11] 93.2 83.1 54.3 59.0 73.3 81.5 18.3
SGPA [ICCV’21] [13] 93.2 88.1 70.7 74.5 82.7 88.4 23.3
GPV-Pose [CVPR’22] [35] 92.9 86.6 67.4 76.2 - 87.4 8.6
Query6DoF [ICCV’23] [15] 91.9 88.1 78.0 83.1 83.9 90.0 19.7
VI-Net [ICCV’23] [18] - - 74.1 81.4 79.3 87.3 28.9
CatFormer [AAAI’24] [12] 93.5 89.9 74.9 79.8 85.3 90.2 -
CLIPose [TCSVT’24] [16] - - 74.8 82.2 82.0 91.2 -
AG-Pose [CVPR’24] [20] 94.1 91.7 77.1 82.0 85.5 91.6 28.9
AG-Pose (DINOv2) [CVPR’24] [20] 94.2 92.5 79.5 83.7 87.1 92.6 11.0
SpherePose [ICLR’25] [21] 94.8 92.4 78.3 84.3 84.8 92.3 -
KeyPose [AAAI’25][36] 94.4 92.6 79.8 83.6 87.1 92.3 -
Ours 94.4 92.9 81.0 85.0 88.1 93.4 16.7
TABLE II: Quantitative comparisons with state-of-the-art methods on the CAMERA25 dataset [10]. The best results are in bold. The symbol indicates that only the trainable parameter count is reported. The symbol -- indicates that the data is not available or not reported.
Figure 4: Qualitative comparisons between INKL-Pose with AG-Pose [20] on the REAL275 dataset, where green and red represent the predicted and ground-truth bounding boxes, respectively. Our INKL-Pose demonstrates superior precision in pose estimation.

### v4

@@ III-D Local Keypoint Feature Aggregator
Next, for each keypoint, we construct a local neighborhood in both coordinate and feature spaces using the $K$-nearest neighbor (KNNK\!N\!N) algorithm. This yields coordinate neighborhoods $\mathcal{P}_{K}\in\mathbb{R}^{N_{kpt}\times K\times 3}$ and feature neighborhoods $\mathcal{F}_{K}\in\mathbb{R}^{N_{kpt}\times K\times d}$, where $K$ is the neighborhood size. We then concatenate the spatial and feature descriptors and apply standard normalization to obtain fused local group features $\mathcal{F}_{K}^{\prime}\in\mathbb{R}^{N_{kpt}\times K\times(d+3)}$, which jointly encode geometric structure and semantic information:
@@
To further enhance the representational capacity of each keypoint, we concatenate the normalized neighbor features $\mathcal{F}^{\prime}_{K}$ with the corresponding central keypoint feature $\mathcal{F}_{kpt}$, resulting in expanded representations $\mathcal{F}^{\prime\prime}_{K}\in\mathbb{R}^{N_{kpt}\times K\times(2\times d+3)}$. This operation allows each keypoint to incorporate both its own feature and contextual information from its spatial neighborhood. We then apply an affine transformation to $\mathcal{F}^{\prime\prime}_{K}$, parameterized by learnable weights $\alpha$ and $\beta$, to capture potential rigid transformations in local geometry and adaptively refine the feature distribution. The transformed feature set $\hat{\mathcal{F}}_{K}\in\mathbb{R}^{N_{kpt}\times K\times(2\times d+3)}$ is computed as follows:
@@ III-E Global Keypoint Feature Aggregator
As illustrated in Fig. 3(b), GKFA employs a bidirectional Mamba-based architecture [gu2024mamba] to efficiently model long-range dependencies and capture object-level structural information. Compared to traditional Transformers, which suffer from quadratic complexity with respect to sequence length, Mamba operates in linear time, making it a more scalable and efficient choice for processing keypoint sequences in point clouds.
To fully exploit Mamba’s bidirectional modeling capabilities, it is necessary to construct both forward and backward feature sequences. A common practice in prior works is to reverse the input sequence order, a technique we refer to as Point Sequence Flipping (PSF). However, point cloud data lack an inherent spatial order, and reversing the index sequence may place geometrically unrelated keypoints next to each other, thereby disrupting the continuity of feature learning. To address this issue, we propose a simple yet effective strategy termed Feature Sequence Flipping (FSF). Rather than reordering keypoints, FSF reverses the feature channels within each keypoint vector while preserving the original spatial arrangement of the point cloud. Formally, given a keypoint feature vector $\mathcal{F}_{local}(i)\in\mathbb{R}^{d}$ expressed as $f_{1}(i),f_{2}(i),\ldots,f_{d}(i)]$, we define FSF as:
$\text{FSF}\left(\mathcal{F}_{\text{local}}\left(i\right)\right)=\left[f_{d}(i),f_{d-1}(i),\ldots,f_{1}(i)\right].$ (9)
@@
$\mathcal{F}_{global}=\text{SSM}\left(\mathcal{F}_{local},\text{FSF}\left(\mathcal{F}_{local}\right)\right)+\mathcal{F}_{local}.$ (10)
@@
To ensure that the aggregated keypoints retain meaningful and expressive geometric structure, we utilize the reconstruction-based supervision mechanism [lin2024instance]. Specifically, we employ an MLP layer to regress dense positional offsets from the global keypoint features. Given $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$, we repeat each feature vector $m=\lfloor N_{rec}/N_{kpt}\rfloor$ times to construct an expanded representation $\mathcal{F}_{global}^{\prime}\in\mathbb{R}^{N_{rec}\times d}$, where $N_{rec}$ is the predefined number of reconstruction points (set to 960 in our experiments). The MLP predicts a set of 3D offsets $\mathcal{O}_{\mathcal{P}}\in\mathbb{R}^{N_{rec}\times 3}$:
@@ III-F Pose and Size Estimator
After obtaining semantically consistent and geometrically informative keypoints and their associated features, we adopt a standard Pose and Size Estimator, as in previous work [lin2022category, lin2024instance], to infer the object’s 6D pose and size. Specifically, we first transform the predicted keypoint coordinates into the NOCS space using the ground-truth rotation $\textit{R}_{gt}$, translation $\textit{t}_{gt}$, and scale $\textit{s}_{gt}$:
@@
We then predict the NOCS keypoints from the extracted features:
@@
$\text{Smooth-}L_{1}(x)=\begin{cases}0.5x^{2},&\text{if }|x|<1\\ |x|-0.5,&\text{otherwise}.\end{cases}$ (16)
@@ IV-A1 Datasets
Following previous works [lin2024instance, ren2025learning], we evaluate the effectiveness of the proposed INKL-Pose on three widely used benchmarks: CAMERA25 [wang2019normalized], REAL275 [wang2019normalized], and the more challenging HouseCat6D [jung2024housecat6d] dataset. CAMERA25 is a large-scale synthetic dataset covering six object categories, containing 300k synthetic RGB-D images with objects rendered on virtual backgrounds, of which 25k images are used for evaluation. REAL275 serves as a real-world counterpart to CAMERA25, covering the same six categories. The training set comprises 4.3k images captured from 7 scenes, while the test set includes 2.75k images from 6 different scenes, with three distinct object instances per category. HouseCat6D is a more complex, real-world multimodal dataset featuring 10 object categories. The training set contains 20k RGB-D images collected across 34 indoor scenes, and the test set includes 3k images from 5 additional scenes. Notably, this dataset poses greater challenges due to its marker-free capture process, dense clutter, and severe occlusion.
@@ IV-A2 Evaluation Metrics
Following previous works [lin2024instance, ren2025learning], we evaluate the performance of INKL-Pose using two widely adopted metrics.
@@
  • nn°mmcm. This metric evaluates the 6D pose accuracy by jointly considering rotation and translation errors. A prediction is regarded as correct if the rotation error is less than nn° and the translation error is below $m$ cm. Following standard protocol, we report results under four commonly used thresholds: 5°2cm, 5°5cm, 10°2cm, and 10°5cm across all three datasets.
@@ IV-A3 Implementation Details
We utilize the off-the-shelf Mask R-CNN [he2017mask] to generate instance masks and resize all input images to 224 $\times$ 224 pixels. For each corresponding object instance, we uniformly sample 1,024 points from the segmented point cloud. To enhance robustness, we apply data augmentation strategies consistent with prior works [lin2024instance, lin2022category], including random rotations (within 0 – 20 degrees), translations (in the range of –0.02 to 0.02 meters), and scalings (from 0.8 to 1.2). For network configuration, the feature dimensions are set as follows: $d_{1}$ = 128, $d_{2}$ = 128 , $d_{3}$ = 64 , and the fused feature dimension is $d=256$. In the local feature aggregation stage, the neighborhood size $K$ is set to 4. Additionally, the entire feature aggregation module consists of $S$ = 12 stacked stages, each comprising both local and global aggregation components to progressively refine the keypoint features. In the keypoint generation stage, the network predicts 96 keypoints per object instance. For keypoint supervision, we sample $N_{fps}$ = 120 reference points to ensure sufficient surface coverage and spatial alignment with the predicted keypoints. In the separation loss, the number of nearest neighbors is set to $M$ = 2, which provides a lightweight yet effective constraint that mitigates keypoint clustering while preserving flexibility in spatial distribution. In the dense point cloud reconstruction stage, each keypoint is repeated 10 times, yielding $N_{rec}$ = 960 reconstructed points. This setting closely matches the resolution of the input point cloud (1,024 points), providing a stable foundation for reconstruction supervision without incurring excessive computational cost.
Regarding optimization, the weights of the loss terms are set as $\{w_{sep},w_{surf},w_{sim},w_{map},w_{pose}\}=\{10.0,\allowbreak 10.0,\allowbreak 15.0,\allowbreak 2.0,\allowbreak 0.3\}$. The model is trained using the Adam optimizer with a cyclical learning rate schedule [smith2017cyclical], where the learning rate varies between $2\times 10^{-5}$ and $5\times 10^{-4}$. All experiments are conducted on a single NVIDIA RTX 4090D GPU with a batch size of 64. For CAMERA25 and REAL275, we train the model using both synthetic and real-world data with a synthetic-to-real ratio of 3:1. For HouseCat6D, the model is trained exclusively on its corresponding dataset.
TABLE I: Quantitative comparisons with state-of-the-art methods on the REAL275 dataset [wang2019normalized]. The best results are in bold. The symbol indicates that only the trainable parameter count is reported. The symbol - indicates that the data is not available or not reported.
@@
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $^{\circ}2\mathrm{cm} 55$$^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$ (M) $\downarrow$
SPD [ECCV’20] [tian2020shape] 77.3 53.2 19.3 21.4 43.2 54.1 18.3
GPV-Pose [CVPR’22] [di2022gpv] 83.0 64.4 32.0 42.9 55.0 73.3 8.6
DPDN [ECCV’22] [lin2022category] 83.4 76.0 46.0 50.7 70.4 78.4 24.6
IST-Net [ICCV’23] [liu2023net] 82.5 76.6 47.5 53.4 72.1 80.5 26.8
Query6DoF [ICCV’23] [wang2023query6dof] 82.5 76.1 49.0 58.9 68.7 83.0 19.7
VI-Net [ICCV’23] [lin2023vi] - - 50.0 57.6 70.8 82.1 28.9
CatFormer [AAAI’24] [yu2024catformer] 83.1 73.8 47.7 53.7 69.0 79.5 -
SecondPose [CVPR’24] [chen2024secondpose] - - 56.2 63.6 74.7 86.0 38.1
AG-Pose [CVPR’24] [lin2024instance] 83.8 77.6 56.2 62.3 73.4 81.2 28.9
AG-Pose (DINOv2) [CVPR’24] [lin2024instance] 84.1 80.1 57.0 64.6 75.1 84.7 11.0
GCE-Pose [CVPR’25] [li2025gce] 84.1 79.8 57.0 65.1 75.6 86.3 -
SpherePose [ICLR’25] [ren2025learning] 84.0 79.0 58.2 67.4 76.2 88.2 -
KeyPose [AAAI’25] [yu2025keypose] 84.2 80.0 57.7 66.0 78.8 88.0 -
MK-Pose [IROS’25] [yang2025mk] 84.0 80.3 60.8 - 78.0 84.6 180.8
CleanPose [ICCV’25] [lin2025cleanpose] - - 61.7 67.6 78.3 86.3 14.7
Ours (DINOv2) 84.0 80.6 63.2 68.3 80.9 88.0 16.7
TABLE II: Quantitative comparisons with state-of-the-art methods on the CAMERA25 dataset [wang2019normalized]. The best results are in bold. The symbol indicates that only the trainable parameter count is reported. The symbol - indicates that the data is not available or not reported.
@@
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $^{\circ}2\mathrm{cm} 55$$^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$ (M) $\downarrow$
SPD [ECCV’20] [tian2020shape] 93.2 83.1 54.3 59.0 73.3 81.5 18.3
GPV-Pose [CVPR’22] [di2022gpv] 92.9 86.6 67.4 76.2 - 87.4 8.6
Query6DoF [ICCV’23] [wang2023query6dof] 91.9 88.1 78.0 83.1 83.9 90.0 19.7
VI-Net [ICCV’23] [lin2023vi] - - 74.1 81.4 79.3 87.3 28.9
CatFormer [AAAI’24] [yu2024catformer] 93.5 89.9 74.9 79.8 85.3 90.2 -
AG-Pose [CVPR’24] [lin2024instance] 94.1 91.7 77.1 82.0 85.5 91.6 28.9
AG-Pose (DINOv2) [CVPR’24] [lin2024instance] 94.2 92.5 79.5 83.7 87.1 92.6 11.0
SpherePose [ICLR’25] [ren2025learning] 94.8 92.4 78.3 84.3 84.8 92.3 -
KeyPose [AAAI’25] [yu2025keypose] 94.4 92.6 79.8 83.6 87.1 92.3 -
MK-Pose [IROS’25] [yang2025mk] 94.1 92.2 77.9 - 86.1 91.7 180.8
CleanPose [ICCV’25] [lin2025cleanpose] - - 80.3 84.2 87.7 92.7 14.7
Ours (DINOv2) 94.4 92.9 81.0 85.0 88.1 93.4 16.7